neat please
Find the Maclaurin series for the function f(x) = cos(x/2). Express your answer in sigma notation.
Answer #1

Solution Given that, )= (ay f(x) = -sin (wa) f(v) = -1. (1982) f(0) = -) f(x) = - corto) f (M = - 4-max.) (2) {(X) = Ludf(0) = Coso = 1 f(o) --- I sino =0 ) = -4.00 C) = Lino 10 =0 we know that Metaurin series is given by :- . h (at z=(x) an(ti ) - 1992 =0 (60) = - - - - - : لم 79 + 9- = ( لا )60) 999 (08% =-1) a23 =0 را !( 2)